# Manifold Learning via Foliations and Knowledge Transfer

### v1

@@ 3.3 Singular Foliations
A foliation on a manifold $M$ is a partition of $M$ into connected immersed submanifolds, called leaves. A foliation is regular if the leaves have the same dimension, singular otherwise. Notice that the map $x\mapsto\mathrm{dim}(\mathcal{L}_{x})$, which associates to $x\in M$ the dimension of its leaf $\mathcal{L}_{x}$, is lower semi-continuous, that is, the dimensions of the leaves in a neighbourhood of $x$ are greater than or equal to $\mathrm{dim}(\mathcal{L}_{x})$ [22]. Whenever we have an equality, we say that $x$ is regular, otherwise we call $x$ singular. It is important to remark that, adhering to the literature [22], the terminology singular point here refers to a point that has a neighbourhood where the leaves have non constant dimension. We can associate a distribution to a foliation by associating to each point, the tangent space to the leaf at that point. Such distribution has constant rank if and only if the foliation is regular. Frobenius Theorem 3.2 applies only to the case of constant rank distributions, however, there are results extending part of Frobenius theorem to non constant rank distributions. In [17] and [30], the authors give sufficient conditions for integrability in this more general setting (see [22] for correct attributions on statements and proofs).
@@
In the practical applications we are interested in, foliations may also have non smooth points, that is points where the leaf through the point, belonging to the foliation, is not smooth. For example, Fig. 4 shows a Xor network with ReLU non linearity displaying both singular and non smooth points.
(a) GeLU.
(b) ReLU.
Figure 4: The learning foliation defined by the distribution $\mathcal{D}$ (4) for a Xor network.
@@
The singular point of each picture is in the center, while non smooth points occur only for Xor with ReLU non linearity. As previously stated, this is a degenerate case, where the involutivity and integrability of $\mathcal{D}$ is granted automatically by low dimensionality, though for (b) in Fig. 4, we cannot define $\mathcal{D}$ at non smooth points.
Remark 3.1.
@@
Notice that Frobenius Theorem 3.2 and its non constant rank counterparts, apply only in the smoothness hypothesis, while for applications, i.e. the case of ReLU networks, it is necessary to examine also non smooth foliations. We plan to explore the non smooth setting more generally in a forthcoming paper.
@@
Let $N(x)=p(y|x,w)=(p_{1}(y|x,w),\dots,p_{c}(y|x,w))$ denote the output of the neural network classifier and $J_{N}(x)=(\nabla_{x}p_{i}(y|x,w))$ its Jacobian matrix. So $\nabla_{x}p_{i}(y|x,w)$ is the $i$-th row (column) of $J_{N}(x)$. One can see, by the very definition of $\mathcal{D}$ (4), that $\mathrm{rank}\mathcal{D}=\mathrm{rank}J_{N}$. We assume $w$ to be constant, that is we fix our model. Let $P=\mathrm{diagonal}(p_{1}(y|x,w),\dots,p_{c}(y|x,w))$.
From now on we assume $N(x)=\mathrm{Softmax}\circ S(x)$ where $S(x)$ represents the score. Then we have:
$J_{N}(x)=(P-p^{t}p)J_{S}(x)$ (6)
where $p=(p_{1}(y|x,w),\dots,p_{c}(y|x,w))$ and $J_{S}(x)$ is the Jacobian of the score.
To study the drop of rank for $\mathcal{D}$ or equivalently for $J_{N}$, let us first look at the kernel of $\left(P-p^{t}p\right)$.
@@
Observation 3.2.
Lemma 3.4 tells us that the rank of the distribution $\mathcal{D}$ or equivalently of $J_{N}(x)=(P-p^{t}p)J_{S}(x)$ is lower at points in the data space where the probability distribution has higher number of $p_{i}\neq 0$. Clearly the points in the dataset, on which our model is trained, are precisely the points where the empirical probability $p$ is mostly resembling a mass probability distribution. Hence at such points, we will observe empirically an average drop of the values of the eigenvalues of the DIM (whose columns generate $\mathcal{D}$), compared to random points in the data space, as our experiments confirm in Sec. 4. As we shall see, this property characterizes the points in the dataset the model was trained with.
In our hypotheses, since the probability vector $p$ is given by a Softmax function, it cannot have null coordinates. Therefore, Lemma 3.4 states that $\dim\ker(P-p^{t}p)=1$ and that the kernel of $P-p^{t}p$ does not depend on the input $x$. Thus, the drops in rank of $\mathcal{D}$ does not depend on $\ker(P-p^{t}p)$ and can only be caused by $J_{S}(x)$.
@@
Let $\mathcal{O}$ be the set of points in $M=\mathbf{R}^{d}$, admitting a neighbourhood where the score function $S$ is non constant. Then, the set of singular points of $J_{S}$, the Jacobian of $S$, is a subset of:
$\bigcup_{k=1}^{\ell-1}f^{-1}_{k}(Z_{f_{k}(M)})\cap\mathcal{O}=\bigcup_{k=1}^{\ell-1}\bigcup_{i=1}^{\dim f_{k}(M)}\left\{x\mid f_{k}(x)_{i}=0\right\}\cap\mathcal{O}$ (10)
@@
If $x_{0}\in\left\{x\mid f_{k}(x)_{i}=0\right\}\cap\mathcal{O}$, then, even if it means making infinitesimal changes to the network weights, there exists a neighborhood of $x_{0}$ such that $(f_{k})_{i}$ is linear on it. Besides, since $x_{0}\in\mathcal{O}$, again, even if it means making infinitesimal changes to the network weights, then $(f_{k})_{i}$ will not be trivial. Thus $\left\{x\mid f_{k}(x)_{i}=0\right\}\cap\mathcal{O}$ is contained in an hyperplane with dimension $d=\dim M$. ∎
Now we see that singular points occur on a (Lebesgue) measure zero set.
@@
$\mathbf{R}^{d}\ni x\mapsto\mathcal{D}_{x}=\mathrm{span}\{\nabla_{x}p_{i}(y|x,w),\,i=1,\dots c\}$ (13)
where $p$ is an empirical probability given by Sofmax and a score function $S$ consisting of a sequence of linear layers and ReLU activations. Then, its singular points (i.e. points where $\mathcal{D}$ changes its rank) are a closed null subset of $\mathbf{R}^{d}$ contained in the union of $\ell$ hypersurfaces, where $\ell$ is the number of layers.
@@
The singular points of $\mathcal{D}$ coincide, by Lemma 3.4 with the points where $J_{S}$ the Jacobian of $S$ as in (8 changes its rank. By Lemma 3.5 this occurs in the union of $\ell$ hypersurfaces. ∎
We conclude by observing that, as a consequence of the proofs of Lemma 3.5 and Thm 3.6, the singular points of the distribution $\mathcal{D}$ are contained in the non smooth points and such points are contained in a measure zero subset of the data space. Hence, if we restrict ourselves to the open set complementing such measure zero set, we can apply Frobenius Thm to $\mathcal{D}$ to get the learning foliation.
@@ 4 Experiments
We perform our experiments on the following datasets: MNIST[23], Fashion-MNIST[49], KMNIST[50] and EMNIST[51], letters only, that we denote with Letters. We also create a dataset that we call CIFARMNIST: it is the CIFAR10 dataset[52] cropped and transformed to be $28\times 28$ gray-scale pictures.
@@
Our neural network is similar to LeNet, with two convolutional layers, followed by a Maxpool and two linear layers with ReLU activation functions, see Fig. 5. This is slightly more general than our hypotheses in Sec. 3.3. The model is then trained on MNIST, reaching $98\%$ of accuracy.
Figure 5: The structure of our CNN - picture created with [25].
@@
In Fig. 6 we compute the Jacobian of the network and we measure its rank by looking at its singular values for 100 sample points in the data space $\mathbf{R}^{784}$ of the MNIST dataset (we normalize the singular values by dividing them by the largest one). The statistical significance of these experiments is detailed in Fig. 7.
(a) Singular values of $J_{N}$ at 100 data points in MNIST.
(b) Singular values of $J_{N}$ at 100 random points.
Figure 6: Singular values of the Jacobian $J_{N}$ of the network, normalized by the largest one.
We see clearly that on points in the dataset the singular values of the Jacobian $J_{N}$, are smaller, as we remarked in Obs. 3.2, as a consequence of our Lemma 3.4.
@@
Figure 7: DIM eigenvalues sorted by decreasing order evaluated on 250 points for each dataset.
@@
We report DIM’s eigenvalues (logarithmic scale) for different datasets in Fig. 7. The vertical segment for each eigenvalue and each dataset represents the values for 80% of the samples, while the colored area represents the values falling in between first and third quartile. The horizontal line represents the median and the triangle represents the mean. The points in MNIST, the training dataset, are clearly identifiable by looking at the colored area.
(a) Validation loss.
(b) Validation accuracy
@@
Figure 8: Loss and accuracy after transfer learning starting from the weights of a ReLU network trained on MNIST ($98\%$ of accuracy) and retraining only the last linear layer.
We perform a proof of concept knowledge transfer by retraining the last linear layer of our model on different datasets. We report in Table 2 the median of highest and lowest DIM eigenvalues in logarithmic scale, their difference $\Delta$ and validation accuracies after retraining of our CNN. We see a correspondence between the median of the lowest eigenvalue and the validation accuracy, suggesting to explore in future works the relation between DIM, foliations and knowledge transfer.
Table 2: Parameters for Knowledge Transfer (logarithmic scale)
Dataset Highest evalue Lowest evalue $\Delta$ DIM Trace Val. Acc.
MNIST -1.78 -8.58 6.70 -1.52 98%
KMNIST 0.49 -7.75 7.76 0.37 75%
Letters 0.11 -7.99 7.82 0.48 80%
Fashion-MNIST 0.14 -8.08 7.76 0.12 81%
CIFARMNIST 0.41 -6.90 6.75 0.27 33%
Noise 0.24 -5.36 5.49 0.27 NA

### v2

@@ 3.3 Singular Foliations and Main Result
A foliation on a manifold $M$ is a partition of $M$ into connected immersed submanifolds, called leaves. A foliation is regular if the leaves have the same dimension, singular otherwise. Notice that the map $x\mapsto\mathrm{dim}(\mathcal{L}_{x})$, which associates to $x\in M$ the dimension of its leaf $\mathcal{L}_{x}$, is lower semi-continuous, that is, the dimensions of the leaves in a neighbourhood of $x$ are greater than or equal to $\mathrm{dim}(\mathcal{L}_{x})$ [lavau2017]. Whenever we have an equality, we say that $x$ is regular, otherwise we call $x$ singular. It is important to remark that, adhering to the literature [lavau2017], the terminology singular point here refers to a point that has a neighbourhood where the leaves have non constant dimension. We can associate a distribution to a foliation by associating to each point, the tangent space to the leaf at that point. Such distribution has constant rank if and only if the foliation is regular. Frobenius theorem˜3.2 applies only to the case of constant rank distributions, however, there are results extending part of Frobenius theorem to non constant rank distributions. In [hermann63] and [nagano66], the authors give sufficient conditions for integrability in this more general setting (see [lavau2017] for correct attributions on statements and proofs).
@@
In the practical applications under consideration, foliations may also have non-smooth points, that is points where the distribution is not continuous, and thus at which the leaf is not smooth. For example, figure˜2 shows a Xor network with ReLU non linearity displaying both singular and non smooth points.
@@
The singular point of each picture is in the center, while non smooth points occur only for Xor with ReLU non linearity. As previously stated, this is a degenerate case, where the involutivity and integrability of $\mathcal{D}$ is granted automatically by low dimensionality, though for (b) in figure˜2, we cannot define $\mathcal{D}$ at non smooth points.
In many interesting applications involving neural networks with ReLU non linearity, the distribution $\mathcal{D}$, generated by the image of the DIM $D$, is both non smooth and singular. It then becomes important to understand where such non smooth and singular points are located. The aim of this section is to show they occupy a measure zero set in the dataspace, so that we can indeed apply successfully Frobenius theorem in a large portion of the dataspace. In our next section we shall investigate experimentally their location in dataspace and the importance in the knowledge transfer questions.
Remark 1.
@@
Notice that Frobenius theorem˜3.2 and its non constant rank counterparts, apply only in the smoothness hypothesis, while for applications, i.e. the case of ReLU networks, it is necessary to examine also non smooth foliations. We plan to explore the non smooth setting more generally in a forthcoming paper.
@@
Let $p:=N(x)=(p(y_{1}|x,w),\dots,p(y_{c}|x,w))$ denote the output of the neural network classifier and $J_{p}(x)=(\nabla_{x}p(y_{i}|x,w))$ its Jacobian matrix. So $\nabla_{x}p(y_{i}|x,w)$ is the $i$-th row (column) of $J_{p}(x)$. In the following, we might omit the $x$ for shorter computations. One can see, by the very definition of $\mathcal{D}$ (6), that $\mathrm{rank}\mathcal{D}=\mathrm{rank}J_{p}$. We assume $w$ to be constant, that is we fix our model. Let $P$ be the diagonal matrix defined by:
$P:=\left(\begin{matrix}p_{1}&&\\ &\ddots&\\ &&p_{C}\end{matrix}\right).$
Recall that $N(x)=p(y|x,w)=\mathrm{Softmax}\circ S(x)$ where $S(x)$ represents the score. Then we have:
$J_{p}(x)=(P-p^{t}p)J_{S}(x)$ (8)
where $J_{S}(x)$ is the Jacobian of the score.
To study the drop of rank for $\mathcal{D}$ or equivalently for $J_{p}$, let us first look at the kernel of $\left(P-p^{t}p\right)$.
@@
Observation 2.
lemma˜3.4 tells us that the rank of the distribution $\mathcal{D}$ or equivalently of $J_{p}(x)=(P-p^{t}p)J_{S}(x)$ is lower at points in the data space where the probability distribution has higher number of $p_{i}=0$. Clearly the points in the dataset, on which our model is trained, are precisely the points where the empirical probability $p$ is mostly resembling a mass probability distribution. Hence at such points, we will observe empirically an average drop of the values of the eigenvalues of the DIM (whose columns generate $\mathcal{D}$), compared to random points in the data space, as our experiments confirm in section˜4. As we shall see, this property characterizes the points in the dataset the model was trained with.
In our hypotheses, since the probability vector $p$ is given by a Softmax function, it cannot have null coordinates. Therefore, lemma˜3.4 states that $\dim\ker(P-p^{t}p)=1$ and that the kernel of $P-p^{t}p$ does not depend on the input $x$. Thus, the drops in rank of $\mathcal{D}$ does not depend on $\ker(P-p^{t}p)$ and can only be caused by $J_{S}(x)$.
@@
The set of singular points of $J_{S}$, the Jacobian of $S$, is a subset of:
$\bigcup_{k=1}^{\ell-1}f^{-1}_{k}(Z_{f_{k}(M)})=\bigcup_{k=1}^{\ell-1}\bigcup_{i=1}^{\dim f_{k}(M)}\left\{x\mid f_{k}(x)_{i}=0\right\}$ (12)
@@
First, note that $f_{k}$ are piecewise affine functions, i.e. divided into a finite number of region on which $f_{k}$ is affine. Let us focus on one of these regions. If $(f_{k})_{i}$ is non-constant, then the equation $f_{k}(x)_{i}=0$ defines a hyperplane of dimension $d=\dim M$ as the kernel of a non-constant affine function. If $(f_{k})_{i}$ is constant, it is either equal to a non-zero value and then the set $\{x\mid f_{k}(x)_{i}=0\}$ is empty; or $(f_{k})_{i}$ is constant equal to zero. This last case may be problematic as it implies that a whole region (with positive measure) could consist of degenerate and non-smooth points. However, note that this only occurs for a specific value of the bias $(b_{k})_{i}$. This is occurring with probability zero after training [^4] To elaborate on this argument, even if this situation were to occur after training, one could escape from it by adding $\varepsilon>0$ to $(b_{k})_{i}$, thereby breaking the problematic equality. With a sufficiently small $\varepsilon$, the network’s predictions should remain close enough to the original ones by continuity., therefore we can safely discard this case. In conclusion, the set of singular points of $J_{S}$ is contained in a union of hyperplanes with dimension $d=\dim M$. ∎
Now we see that singular points occur on a (Lebesgue) measure zero set given by the union of the hypersurfaces of equation $(f_{k})_{i}=0$, which is a piecewise linear condition (see proof of lemma˜3.5).
@@
$\mathbf{R}^{d}\ni x\mapsto\mathcal{D}_{x}=\mathrm{span}\{\nabla_{x}p(y_{i}|x,w),\,i=1,\dots c\}$ (15)
where $p$ is an empirical probability given by Softmax and a score function $S$ consisting of a sequence of linear layers and ReLU activations. Then, its singular points (i.e. points where $\mathcal{D}$ changes its rank) are a closed null subset of $\mathbf{R}^{d}$ contained in a union hypersurfaces.
@@
The singular points of $\mathcal{D}$ coincide, by lemma˜3.4 with the points where $J_{S}$ the Jacobian of $S$ as in (10) changes its rank. By lemma˜3.5 this occurs in a union of hypersurfaces, given by the equations $(f_{k})_{i}=0$ (see proof of lemma˜3.5). ∎
As a consequence of the proofs of lemma˜3.5 and theorem˜3.6, the singular points of the distribution $\mathcal{D}$ are contained in the non smooth points and such points are contained in a measure zero subset of the data space. Hence, if we restrict ourselves to the open set complementing such measure zero set, we can apply Frobenius Theorem 3.2 to the distribution $\mathcal{D}$ to get a foliation. Since it is of significance in the dataspace, as we are going to elucidate in the next section, we call such foliation the data foliation.
Remark 2.
theorem˜3.6 holds also in the more general hypothesis of piecewise linear activation functions, in place of ReLU.
Observation 3.
There is empirical evidence [Hanin2019ComplexityOL] that, removing the measure zero set of non smooth points as in theorem˜3.6, partitions the dataspace into disjoint connected components. In each component the distribution $\mathcal{D}$ has constant rank, hence the foliation restricted to one component is regular. However the global foliation on the dataspace, granted locally by Frobenius Theorem 3.2, may vary its rank as we move from one connected component to another, as the experiments in our next section suggest. Hence the notion of singular foliation, i.e. a foliation with leaves of different dimensions, become essential to properly model the dataspace, in which we train the neural network.
Interpretation of Singularities.
The dimension of a leaf from the data foliation indicates the number of degrees of freedom available to modify the output prediction by adjusting the input. When the rank of $\mathcal{D}_{x}$ (or equivalently $D(x)$) is lower than c1c-1, then the changes to the prediction are constrained. For instance, if the rank of $\mathcal{D}_{x}$ is only $1$, then any local change to the input will produce a displacement along a one dimensional path in the output space. A singular point is then where leaves of different dimensions meet in the input space, adding (or removing in the opposite direction) degrees of freedom to the output’s modifications.
@@ 4 Experiments
In this section we look at some experiments to elucidate the role of the singular and non-smooth points (see theorem˜3.6) for the distribution $\mathcal{D}$, generated by the image of the DIM $D$ (see Def. (4)), in knowledge transfer questions. The emergence of a singular foliation structure in an dense open set of the dataspace, granted by theorem˜3.6, via Frobenius theorem˜3.2 (see also ˜3) is related with the inner structure of datasets in such dataspace. Indeed, we will show that an effective way to measure the "distance" between different datasets embedded in the same dataspace, is related with the rank of $\mathcal{D}$, or equivalently $D$, and the (relative) magnitude of $D$’s eigenvalues.
@@
We perform our experiments on the following datasets: MNIST [lecun1998], Fashion-MNIST [fashionmnist], KMNIST [kmnist] and EMNIST [emnist], letters only, that we denote with Letters. We also create a dataset that we call CIFARMNIST: it is the CIFAR10 dataset [cifar] cropped and transformed to be $28\times 28$ gray-scale pictures. Finally, we create a randomly generated dataset called Noise as a baseline, with input points uniformly sampled in the $0,1]^{28\times 28}$ cube.
@@
Our neural network is similar to LeNet, with two convolutional layers, followed by an Averagepool and two linear layers with ReLU activation functions , see Appendix A for more details on the architecture. This is slightly more general than our hypotheses in section˜3.3. The model is then trained on MNIST, reaching $98\%$ of accuracy.  [^5] All the code used in this section is available at https://github.com/eliot-tron/SingFolDIM.
@@
In figure˜3 we compute the DIM and we measure its rank by looking at its eigenvalues for 100 sample points in the data space $\mathbf{R}^{784}$ of the MNIST dataset and on 100 uniformly random points in $0,1]^{784}$. The statistical significance of these experiments is detailed in figure˜4.
(a) MNIST data points.
(b) Uniformly random points (Noise).
@@
Figure 3: DIM eigenvalues sorted by descending order evaluated on 100 points. Each blue line correspond to one data point.
We see clearly that on points in the training set the eigenvalues of the DIM $D$, are smaller. Since $\mathrm{rank}D=\mathrm{rank}\mathcal{D}$ at each point (see section˜3.2), the drop of the eigenvalues of $J_{p}$ reflects the change in rank of the distribution $\mathcal{D}$ in the proximity of datapoints, i.e. the points on which network was trained. This provides evidence for modeling the dataspace via a singular foliation (see ˜3. and lemma˜3.4).
Figure 4: DIM eigenvalues sorted by descending order, evaluated on 10,000 points for each dataset, for a network trained on MNIST.
We compute the eigenvalues $\left(\lambda_{i}\right)_{1\leq i\leq d}$ of the DIM and sort them as such:
$\lambda_{(1)}\leq\ldots\leq\lambda_{(d)}.$
Then, we report the eigenvalues $\lambda_{(d)},\ldots,\lambda_{(d-9)}$, sorted in descending order [^6] Eigenvalues $\lambda_{(d-10)},\ldots,\lambda_{(1)}$ are not plotted because $\mathrm{rank}D<c=10$, and therefore are equal to zero. $\lambda_{(d-9)}$ is plotted only for reference of what the numerical zero is. (logarithmic scale), for different datasets in Fig. 4.
@@
The vertical segment for each eigenvalue and each dataset represents the values for 80% of the samples, while the colored area represents the values falling in between first and third quartile. The horizontal line represents the median and the triangle represents the mean. The points in MNIST, the training dataset, are clearly identifiable by looking at the colored area. On the opposite, the randomly generated points in Noise yield higher eigenvalues with fewer variations. This difference in eigenvalues between the various datasets prompts us to consider using the DIM to measure how different the datasets are from the network’s perspective.
To provide a point of comparison to our analysis, we have included with figure˜5 a similar plot to figure˜4 but for a neural network with random weights (i.e. before any training). As can be seen, the singularities appear during training when they become concentrated around points similar to those in the training set.
Figure 5: DIM eigenvalues sorted by descending order, evaluated on 10,000 points for each dataset, for a network with random weights.
@@
Figure 6: Loss and accuracy after knowledge transfer starting from the weights of a ReLU network trained on MNIST ($98\%$ of accuracy) and retraining only the last linear layer.
We perform a proof of concept knowledge transfer experiment by retraining the last linear layer of our model on different datasets. This experiment was selected as it relies on the degree of similarity between datasets from the perspective of the model. We report in figure˜6 the evolution of the loss (dotted) and accuracy (plain) of the network after 30 epochs of retraining on 4 different datasets. We notice a correspondence between the magnitude of the eigenvalues and the validation accuracy suggesting to explore in future works the relationship between DIM, foliations and knowledge transfer. Indeed, CIFARMNIST performs worst than the others and has higher eigenvalues in average.
With the magnitude of the eigenvalues of the DIM, we find similar results as the authors of [Alvarez-Melis2019Geometric] regarding the similarity between MNIST and the other datasets.
We report in table˜2 the median of highest ($\lambda_{(d)}$) and lowest non-zero ($\lambda_{(d-8)}$) DIM eigenvalues in logarithmic scale, $\Delta$ the average length of the vertical segments in figure˜4, and the validation accuracy after retraining of our CNN. We see a correspondence between the median of the lowest non-zero eigenvalue and the validation accuracy, suggesting to explore in future works the relation between DIM, foliations and knowledge transfer. The significance of the last non-zero eigenvalue of the DIM lies in its correspondence with the rank of the foliation. Indeed, the eigenvalue $\lambda_{(d-8)}$ being (close to) zero means that the rank drops at (close to) the data point. This supports our conjecture that similar datasets (i.e. with high knowledge transfer potential) will lead to data leaves with similar geometric properties. Moreover it suggests a definition of dataset distance based on such eigenvalue magnitude.
Table 2: Parameters for Knowledge Transfer (log scale)
Dataset $\lambda_{(d)}$ $\lambda_{(d-8)}$ $\Delta$ DIM Trace Val. Acc.
MNIST -3.10 -13.40 10.07 -1.52 98%
Fashion-MNIST 0.57 -11.54 11.56 0.12 81%
Letters 0.09 -12.56 11.99 0.48 80%
KMNIST 0.34 -11.17 10.98 0.37 75%
CIFARMNIST 0.67 -9.34 9.44 0.27 33%
Noise 1.17 -6.66 7.58 0.27 NA